You have been told that you need $15,000 today for every $50,000 you want when you retire 30 years from now. What rate of interest was used in the present value computation? Assume interest is compounded annually
Answer #1

This question requires application of basic time value of money function, according to which

FV = PV * (1 + r)n

$50,000 = $15,000 * (1 + r)30

3.3333 = (1 + r)30

1.0409 = 1 + r

r = 0.0409

r = 4.09% --> Answer